# Parity erasure: a foundational principle for indefinite causal order

### v1

@@ Abstract
The process matrix formalism is built on the concept of higher-order quantum operations, also known as quantum supermaps, which generalizes the quantum circuit model [14, 15]. Although supermaps can be defined within arbitrary general operational probabilistic theory (OPT) [1, 25, 26, 27, 6, 28, 29, 30, 31, 32] admitting the causality principle [27], processes with ICO remain largely abstract mathematical constructs lacking a clear operational interpretation. The possibility of ICO therefore opens a new frontier for the operational approach: if nature indeed allows such processes, it becomes crucial to understand whether the extension to ICO can be derived from operational principles without assumptions for higher-order operations.
In this paper, we identify a universal property of supermaps that holds in every causal OPT. Specifically, we show that supermaps satisfy a set of input–output constraints that we term parity erasure. Moreover, we prove that the parity erasure principle fully characterizes supermaps in every causal OPT which additionally satisfies the principle of local tomography and a property termed one-way-signaling decomposability. In particular, the process matrix formalism, which characterizes quantum supermaps, can be derived from the parity erasure principle. Our results thus provide an information-theoretic axiomatization of the process matrix formalism, revealing a fundamental constraint on information exchange within an indefinite causal structure.
@@
Following the construction of the process matrix formalism [15], we consider a general OPT scenario [16] where there are $N$ separate experimenters, each of which can freely choose an operation in their laboratory. The operation of the $i$-th experimenter, denoted by $\left(\mathcal{M}_{a_{i}}^{(i)}\right)$, is a test consisting of probabilistic transformations from system $A_{i}$ to $X_{i}$ and producing an outcome $a_{i}$. The environment outside their laboratories is described by a supermap $\mathcal{S}$ transforming the local operations such that there is a valid joint probability distribution $p\left(\mathcal{M}_{a_{1}}^{(1)},\mathcal{M}_{a_{2}}^{(2)},\cdots\mid\mathcal{S}\right)$ for every choice of $\left(\mathcal{M}^{(i)}\right)_{i=1}^{N}$. Moreover, the map $\mathcal{S}$ is completely positive, that is, for every transformation $\widetilde{\mathcal{M}}^{(i)}$ from the composite system $A_{i}\otimes A^{\prime}_{i}$ to the composite system $X_{i}\otimes X^{\prime}_{i}$, the map $\mathcal{S}\otimes\mathcal{I}^{{\mathsf{Transf}}(A^{\prime}\to X^{\prime})}$ produces a valid transformation from the composite system $A^{\prime}:=\bigotimes_{i=1}^{N}A^{\prime}_{i}$ to the composite system $X^{\prime}:=\bigotimes_{i=1}^{N}X^{\prime}_{i}$, where $\mathcal{I}^{{\mathsf{Transf}}(A^{\prime}\to X^{\prime})}$ denotes the identity supermap on the set of transformations ${\mathsf{Transf}}(A^{\prime}\to X^{\prime})$. See Fig. 1(a) for an illustration in the case of two parties.
In addition, we assume the physicalization of readout [33], that is, every test can be realized as a channel followed by a measurement on an additional system, which indicates that supermaps can be completely characterized by their behaviors on channels.
@@
The causality principle imposes a basic constraint on the local input-output relations of supermaps, which forbids signaling from the future to the past in experiments in local laboratories. A similar constraint is studied in Ref. [34] in a framework beyond supermaps. Therefore, for every party $i$, the marginal probability distribution $p(a_{i}\mid x_{1},\cdots,x_{N})$ is independent of the value of $x_{i}$.
@@
In the two-party case, classical supermaps are known to be causally separable [15, 35], that is, every two-party classical supermap can be decomposed as a random mixture of ones compatible with fixed causal orders. When there is a fixed causal order between the two parties, one of the bits $x_{1}$ and $x_{2}$ cannot be decoded from the joint outcome $(a_{1},a_{2})$ in any local input-output relation, as causally-ordered classical supermaps can be realized by a circuit that discards the bit of the last party. Therefore, with a fixed causal order between the two parties or a random mixture of causal orders, the information about the parity $x_{1}\oplus x_{2}$ is not available from the joint measurement outcome $(a_{1},a_{2})$. It follows that the parity $x_{1}\oplus x_{2}$ cannot be predicted from the local input-output relations of any two-party supermap. This parity erasure property can be generalized to an arbitrary number of parties, which is formally stated in the following theorem.
@@
Figure 2: One-way-signaling decomposability. Under this assumption, every transformation satisfying the no-influence condition $X_{1}\not\to A_{1}$ can be implemented by a circuit with a slot from $A_{1}$ to $X_{1}$.
@@
Every bipartite channel $\mathcal{U}\in{\mathsf{Transf}}_{1}(X_{1}\otimes X_{2}\to A_{1}\otimes A_{2})$ that satisfies the no-influence condition $X_{1}\not\to A_{1}$ (that is, the marginal output on the subsystem $A_{1}$ does not depend on the marginal input on the subsystem $X_{1}$) can be implemented by concatenating a transformation $\mathcal{E}\in{\mathsf{Transf}}_{1}(X_{2}\to A_{1}\otimes M)$ and another transformation $\mathcal{D}\in{\mathsf{Transf}}_{1}(X_{1}\otimes M\to A_{2})$ through a system $M$, as visualized in Fig. 2.
For classical channels, one-way-signaling decomposability follows directly from the chain rule of probability theory, and this property has also been established for quantum channels [37]. This property guarantees that any channel for which a given input subsystem cannot influence a given output subsystem admits an explicit circuit realization in which an arbitrary transformation between the two subsystems can be probed.
With local tomography and one-way-signaling decomposability, it follows that any $N$-partite parity-erasure channel can act as a supermap through a procedure in which the local transformations of the $N$ parties are inserted one-by-one into a sequence of decompositions of the parity-erasure channel. The proof is established by a recursive property of extended parity-erasure channels. Extended $k$-partite parity-erasure channels from $X_{1}\otimes\cdots\otimes X_{k}\otimes P$ to $A_{1}\otimes\cdots\otimes A_{k}\otimes F$ refer to those that are reduced to parity-erasure channels after preparing an arbitrary state on the system $P$ and discarding the system $F$. Such channels admit a one-way-signaling decomposition with a slot from $A_{i}$ to $X_{i}$ ($i\in\{1,\cdots,k\}$) because the parity-erasure property for a single party reduces to the no-influence condition $X_{i}\not\to A_{i}$. The recursive property is stated in the following lemma (see the proof in Appendix B).
@@
Discussion and outlook.— We have provided an information-theoretic characterization of processes with ICO by identifying the parity erasure principle that is universally satisfied by supermaps in all causal OPTs. Under the minimal assumptions of local tomography and one-way-signaling decomposability, this principle determines the admissible set of supermaps and, when applied to quantum theory, yields precisely the process matrix formalism. Our results place ICO within the operational reconstruction program and show that it can be seen as arising from a fundamental property of information exchange rather than from purely formal extensions of the circuit model.
@@
Acknowledgment.— This work was supported by the F.R.S.-FNRS under project CHEQS within the Excellence of Science (EOS) program, by the ID# 62312 grant from the John Templeton Foundation, as part of the “The Quantum Information Structure of Spacetime” Project (QISS), and by the ID# 63683 grant from the John Templeton Foundation, as part of the “WithOut SpaceTime” Project (WOST). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. O. O. is a Senior Research Associate of the Fonds de la Recherche Scientifique (F.R.S.–FNRS).
@@ Appendix A Proof of Theorem 1
where we have used the notation $_{[X]}S:=\operatorname{Tr}_{X}[S]\otimes\frac{I_{X}}{d_{X}}$ for a quantum system $X$ of dimension $d_{X}$, and the shorthand $_{[\sum\alpha_{X}X]}S:=\sum_{X}\alpha_{X}\cdot{}_{[X]}S$ for a linear combination of the terms $\{{}_{[X]}S\}_{X}$ associated with a collection of systems indexed by $X$ (for example, $_{[(1-X)Y]}S$ serves as a shorthand for $_{[Y]}S-{}_{[XY]}S$).
@@
Conditions (2) and (3) are necessary for $S$ to represent a quantum channel. Conditions (4) are equivalent to the following equations:

### v2

@@ Abstract
The process matrix formalism is built on the concept of higher-order quantum operations, also known as quantum supermaps, which generalizes the quantum circuit model [14, 15]. Although supermaps can be defined within any operational probabilistic theory (OPT) [1, 25, 26, 27, 6, 28, 29, 30, 31, 32] admitting the causality principle [27], processes with ICO remain largely abstract mathematical constructs lacking a clear operational interpretation. The possibility of ICO therefore opens a new frontier for the operational approach: if nature indeed allows such processes, it becomes crucial to understand whether the extension to ICO can be derived from operational principles without assumptions for higher-order operations.
In this paper, we identify a universal property of supermaps that holds in every causal OPT. Specifically, we show that supermaps satisfy a set of input–output constraints that we term parity erasure. Moreover, we prove that the parity erasure principle fully characterizes supermaps in every causal OPT which additionally satisfies the principle of local tomography and a property termed one-way-signaling decomposability. In particular, the process matrix formalism, which describes quantum supermaps, can be derived from the parity erasure principle. Our results thus provide an information-theoretic axiomatization of the process matrix formalism, revealing a fundamental constraint on information exchange within an indefinite causal structure.
@@
Following the construction of the process matrix formalism [15], we consider a general OPT scenario [16] where there are $N$ separate experimenters, each of which can freely choose an operation in their laboratory. The operation of the $i$-th experimenter, denoted by $\left(\mathcal{M}_{a_{i}}^{(i)}\right)$, is a test consisting of transformations from a system $A_{i}$ to a system $X_{i}$ with an outcome $a_{i}$. The environment outside their laboratories is described by a supermap $\mathcal{S}$ transforming the local operations such that there is a valid joint probability distribution $p\left(\mathcal{M}_{a_{1}}^{(1)},\mathcal{M}_{a_{2}}^{(2)},\cdots\mid\mathcal{S}\right)$ for every choice of $\left(\mathcal{M}^{(i)}\right)_{i=1}^{N}$. Moreover, the map $\mathcal{S}$ is completely positive, that is, for every transformation $\widetilde{\mathcal{M}}^{(i)}$ from the composite system $A_{i}\otimes A^{\prime}_{i}$ to the composite system $X_{i}\otimes X^{\prime}_{i}$, the map $\mathcal{S}\otimes\mathcal{I}^{{\mathsf{Transf}}(A^{\prime}\to X^{\prime})}$ produces a valid transformation from the composite system $A^{\prime}:=\bigotimes_{i=1}^{N}A^{\prime}_{i}$ to the composite system $X^{\prime}:=\bigotimes_{i=1}^{N}X^{\prime}_{i}$, where $\mathcal{I}^{{\mathsf{Transf}}(A^{\prime}\to X^{\prime})}$ denotes the identity supermap on the set of transformations ${\mathsf{Transf}}(A^{\prime}\to X^{\prime})$. See Fig. 1(a) for an illustration in the case of two parties.
In addition, we assume the physicalization of readout [33], that is, every test can be realized as a channel followed by a measurement on an auxiliary system, which indicates that supermaps can be defined by their behaviors on channels.
@@
The causality principle imposes a basic constraint on the local input-output relations of supermaps, which forbids signaling from the future to the past in experiments in local laboratories. A similar constraint is studied in Refs. [34, 35] in a framework beyond supermaps. Therefore, for every party $i$, the marginal probability distribution $p(a_{i}\mid x_{1},\cdots,x_{N})$ is independent of the value of $x_{i}$.
@@
In the two-party case, classical supermaps are known to be causally separable [15, 36], that is, every two-party classical supermap can be decomposed as a random mixture of ones compatible with fixed causal orders. When there is a fixed causal order between the two parties, one of the bits $x_{1}$ and $x_{2}$ cannot be decoded from the joint outcome $(a_{1},a_{2})$ in any local input-output relation, as causally ordered classical supermaps can be realized by a circuit that discards the bit of the last party. Therefore, with a fixed causal order between the two parties or a random mixture of causal orders, the information about the parity $x_{1}\oplus x_{2}$ is not available from the joint measurement outcome $(a_{1},a_{2})$. It follows that the parity $x_{1}\oplus x_{2}$ cannot be predicted from the local input-output relations of any two-party supermap. This parity erasure property can be generalized to an arbitrary number of parties, which is formally stated in the following theorem.
@@
Figure 2: One-way-signaling decomposability. Under this assumption, every channel satisfying the no-influence condition $X_{1}\not\to A_{1}$ can be implemented by a circuit with a slot from $A_{1}$ to $X_{1}$.
@@
Every bipartite channel from $X_{1}\otimes X_{2}$ to $A_{1}\otimes A_{2}$ that satisfies the no-influence condition $X_{1}\not\to A_{1}$ (that is, the marginal output of the system $A_{1}$ does not depend on the input on the system $X_{1}$) can be implemented by connecting a channel $\mathcal{E}\in{\mathsf{Transf}}_{1}(X_{2}\to A_{1}\otimes M)$ and a channel $\mathcal{D}\in{\mathsf{Transf}}_{1}(X_{1}\otimes M\to A_{2})$ through a memory system $M$, as visualized in Fig. 2.
For classical channels, one-way signaling decomposability follows directly from the chain rule of probability theory, and this property has been established for quantum channels [38] and, more generally, for channels in any causal OPT with local tomography and purification [27]. This guarantees that any channel in which a given input subsystem cannot influence a given output subsystem admits an explicit circuit realization, where an arbitrary transformation acting between the two subsystems can be probed.
With local tomography and one-way-signaling decomposability, it follows that every $N$-partite parity-erasure channel can act as a supermap through a procedure in which the local transformations of the $N$ parties are inserted one-by-one into a sequence of decompositions of the parity-erasure channel. The proof is established by a recursive property of extended parity-erasure channels. Extended $k$-partite parity-erasure channels from $X_{1}\otimes\cdots\otimes X_{k}\otimes P$ to $A_{1}\otimes\cdots\otimes A_{k}\otimes F$ refer to those that become parity-erasure channels after preparing an arbitrary state on the system $P$ and discarding the system $F$. Such channels admit a one-way-signaling decomposition with a slot from $A_{i}$ to $X_{i}$ ($i\in\{1,\cdots,k\}$) because the parity-erasure property for a single party reduces to the no-influence condition $X_{i}\not\to A_{i}$. The recursive property is stated in the following lemma (see the proof in Appendix B).
@@
Discussion and outlook.— We have provided an information-theoretic axiomatization of processes with ICO by identifying the parity erasure principle that is universally satisfied by supermaps in all causal OPTs. Under the assumptions of local tomography and one-way-signaling decomposability, this principle determines the admissible set of supermaps and, when applied to quantum theory, yields precisely the process matrix formalism. Our results place ICO within the operational reconstruction program and show that it can be seen as arising from a fundamental property of information exchange rather than from purely formal extensions of the circuit model.
@@
Acknowledgment.— Z. L. is grateful to G. Chiribella for helpful comments. This work was supported by the F.R.S.-FNRS under project CHEQS within the Excellence of Science (EOS) program, by the ID# 62312 grant from the John Templeton Foundation, as part of the “The Quantum Information Structure of Spacetime” Project (QISS), and by the ID# 63683 grant from the John Templeton Foundation, as part of the “WithOut SpaceTime” Project (WOST). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. O. O. is a Senior Research Associate of the Fonds de la Recherche Scientifique (F.R.S.–FNRS).
@@ Appendix A Proof of Theorem 1
where we have used the notation $_{[X]}S:=\operatorname{Tr}_{X}[S]\otimes\frac{\mathbb{I}_{X}}{d_{X}}$ for a quantum system $X$ of dimension $d_{X}$, and the shorthand $_{[\sum\alpha_{X}X]}S:=\sum_{X}\alpha_{X}\cdot{}_{[X]}S$ for a linear combination of the terms $\{{}_{[X]}S\}_{X}$ associated with a collection of systems indexed by $X$ (for example, $_{[(1-X)Y]}S$ serves as a shorthand for $_{[Y]}S-{}_{[XY]}S$).
@@
Conditions (4) are equivalent to the following equations: